OpenStudy (anonymous):

Which point lies on the line defined by 3x + 6 y = 2? PLEASE SHOW WORK :]

OpenStudy (anonymous):

Are there particular points you want us to look at?

OpenStudy (anonymous):

An infinite number of points lie on a line

OpenStudy (anonymous):

If you want any point, you can let x=0 to get 6y=2 y=1/3 you would have the point (1, 1/3) on the line

OpenStudy (anonymous):

one of the answer choices was (1, -1/3)

OpenStudy (anonymous):

err that was a mistake, the point i got was (0, 1/3) but (1, -1/3) might be a point... let's see :)

OpenStudy (anonymous):

if x=1, then 3+6y=2 6y= -1 y= -1/6 So (1, -1/3) would not lie on th eline

OpenStudy (anonymous):

hmmm other answer choices were (0,2), (0,6),(1,-1/6)

OpenStudy (anonymous):

well we have the answer then. We just solved for the point on the line when x=1 and got (1, -1/6) :) Do you understand the process though of how we got the answer?

OpenStudy (anonymous):

I do...I just plugged in all the answers to the first equation (3x+6y=2) and (1,-1/6) was the only answer that worked...thats another way to solve it right?

OpenStudy (anonymous):

yup :) That works too

OpenStudy (anonymous):

Cool ! <3 Thank you !

OpenStudy (anonymous):

Ive got a CST tomorrow...im scared